OpenStudy (anonymous):

@DanJS In which quadrant does the solution of the system fall? a. 1 b. 2 c. 3 d. 4 In which quadrant does the solution of the system fall? a. 1 b. 2 c. 3 d. 4 The cost of four scarves and six hats is $52.00. The cost of two hats is $1.00 more than the cost of one scarf. What is the cost of one scarf? a. $4.00 b. $5.00 c. $6.00 d. $7.00 Please help!

OpenStudy (anonymous):

oops i forgot to put the equations for the first two

OpenStudy (anonymous):

Here: #1: y=x-1 and y=-3x-5 #2: x+y=4 and 2x-y=2

OpenStudy (anonymous):

@DanJS hurry!!!

OpenStudy (danjs):

sorry

OpenStudy (danjs):

y = x - 1 y = -3x -5

OpenStudy (danjs):

Y= something Y = else something = else

OpenStudy (danjs):

solve x - 1 = -3x - 5

OpenStudy (anonymous):

idk how really ;( i sent an email to my teacher early this morning but she hasnt replied:(

OpenStudy (danjs):

ok, i will help you, as long as you know how to add and subtract we are good

OpenStudy (anonymous):

i do duh :P

OpenStudy (danjs):

y = x - 1 y = -3x -5 Here they tell you y = something, and y = something else, y = both of those things, so we can set both of those things equal to each other... x - 1 = -3x - 5 see that part?

OpenStudy (anonymous):

yes i do and i understand it

OpenStudy (danjs):

x - 1 = -3x - 5 The objective is to isolate x all by itself. You can do that by adding, subtracting, multiplying, whatever you want to do. You just have to make sure you do whatever on both sides of the equal sign , so that you dont alter the original equation.

OpenStudy (anonymous):

so what do i do?

OpenStudy (anonymous):

yep ok

OpenStudy (danjs):

The first thing, is to move any of the terms with an X to the same side, you can ADD 3x to both sides of the equation... x + 3x - 1 = -3x + 3x -5

OpenStudy (anonymous):

yep

OpenStudy (danjs):

you can add together terms containing x, x + 3x = 4x, and -3x + 3x = 0

OpenStudy (danjs):

4x - 1 = -5

OpenStudy (anonymous):

oh and then now you add one to both sides, right?

OpenStudy (anonymous):

4x=-4?

OpenStudy (danjs):

yep, you are getting the hang of it

OpenStudy (anonymous):

and then that would be x=-1

OpenStudy (danjs):

exactly, you just divided both sides by 4

OpenStudy (danjs):

Now that you know X = -1 you can use that information in one of the original equations, and find Y y = x - 1 y = -3x -5

OpenStudy (danjs):

Y should be the same if you put -1 in either of those equations

OpenStudy (anonymous):

well actually i dont need to do that. now i graph it and find out what box it is in

OpenStudy (anonymous):

1 2 3 or 4

OpenStudy (anonymous):

do you know? and can you help me graph it so i can figure this out?

OpenStudy (danjs):

you really don't have to graph it

OpenStudy (danjs):

if you found x = -1, and find y from one of the equatons, y = x - 1 ; x = -1 y = -1 - 1 = -2 y=-2 and x=-1 is the solution to those two given equations

OpenStudy (danjs):

that is the point (x,y) = ( -1, -2)

OpenStudy (anonymous):

so it would be in box 3

OpenStudy (danjs):



OpenStudy (danjs):

right

OpenStudy (danjs):

that is the intersection point , if you graphed those two lines

OpenStudy (anonymous):

yep thanks:) now next question please

OpenStudy (danjs):

x + y = 4 2x - y = 2 There are a couple ways to solve this, you could do the same thing as the last one, if you solved both for y = first, then set them equal like last time

OpenStudy (danjs):

Here is the way i would do it...

OpenStudy (danjs):

x + y = 4 2x - y = 2 ADD THEM TOGETHER ------------- 3x + 0y = 6 3x = 6

OpenStudy (danjs):

see how the Y variable is 0 when you add the two equations together

OpenStudy (anonymous):

ok x is 2 then

OpenStudy (anonymous):

so what would be the equation to graph it?

OpenStudy (danjs):

it just asks , what quadrant is the solution in, same as last time, you have to figure what Y is first when x is 2

OpenStudy (danjs):

x + y = 4 2x - y = 2 and x=2 what is y?

OpenStudy (anonymous):

ok well we know x is 2, so the equation would be: i think it would be 0 cause what i got is 0y=-2x

OpenStudy (danjs):

where did you get 0y = -2x from? Use one of the original equations, and put in x=2, x + y = 4 2x - y = 2

OpenStudy (anonymous):

oh ok. it would be y+2x=2 and then you subtract 2x from both sides, so it would be y=-2x+2 i think :P :(

OpenStudy (danjs):

no, replace x with 2 x + y = 4 2 + y = 4 or 2x - y = 2 2*2 - y = 2

OpenStudy (anonymous):

so 4-y=2

OpenStudy (anonymous):

so y is 2 because 4-2=2

OpenStudy (danjs):

yes, 2 + y = 4 all you have to do to get y, is take off 2 from both sides y = 4-2 = 2

OpenStudy (anonymous):

so x is 2 and y is 2 it would be in box 2

OpenStudy (anonymous):

right?

OpenStudy (danjs):